OpenStudy (anonymous):

The length of human pregnancies from conception to birth varies according to a distribution that is approximately normal with mean 266 days and standard deviation 16 days. How long do the longest 20% of pregnancies last?

OpenStudy (dumbcow):

using a standard normal table, obtain the z-value with 80% to left/ 20% to right http://en.wikipedia.org/wiki/Z_table#Cumulative_table use "z-score" formula to obtain num of days(N) \[Z = \frac{N-mean}{std dev}\] \[0.84 = \frac{N-266}{16}\] \[N = 279.44\] -The longest 20% last more than 279 days